### v1

# A Case Study of Cross-Lingual Zero-Shot Generalization for Classical Languages in LLMs

Abstract
Large Language Models (LLMs) have demonstrated remarkable generalization capabilities across diverse tasks and languages. In this study, we focus on natural language understanding in three classical languages—Sanskrit, Ancient Greek and Latin—to investigate the factors affecting cross-lingual zero-shot generalization. First, we explore named entity recognition and machine translation into English. While LLMs perform equal to or better than fine-tuned baselines on out-of-domain data, smaller models often struggle, especially with niche or abstract entity types. In addition, we concentrate on Sanskrit by presenting a factoid question–answering (QA) dataset and show that incorporating context via retrieval-augmented generation approach significantly boosts performance. In contrast, we observe pronounced performance drops for smaller LLMs across these QA tasks. These results suggest model scale as an important factor influencing cross-lingual generalization. Assuming that models used such as GPT-4o and Llama-3.1 are not instruction fine-tuned on classical languages, our findings provide insights into how LLMs may generalize on these languages and their consequent utility in classical studies.
A Case Study of Cross-Lingual Zero-Shot Generalization for
Classical Languages in LLMs
V.S.D.S. Mahesh Akavarapuα, Hrishikesh Terdalkarβ, Pramit Bhattacharyyaγ, Shubhangi Agarwalβ, Vishakha Deulgaonkarγ, Pralay Mannaγ, Chaitali Dangarikarγ, Arnab Bhattacharyaγ αUniversity of Tübingen, βUniversity of Lyon 1, γIndian Institute of Technology Kanpur mahesh.akavarapu@uni-tuebingen.de, hrishikesh.terdalkar@liris.cnrs.fr, arnabb@cse.iitk.ac.in
1 Introduction
Large Language Models (LLMs) (Brown, 2020; Ouyang et al., 2022; Touvron et al., 2023) are known to generalize across various tasks using data from languages present in their pre-training phase, even when not present in instruction tuning (Wang et al., 2022; Muennighoff et al., 2023; Han et al., 2024). Previous work has demonstrated generalization to several low-resource languages via few-shot in-context learning (Cahyawijaya et al., 2024). In this study, we focus on zero-shot generalization to natural language understanding (NLU) tasks for three Indo-European classical languages—Sanskrit (san), Ancient Greek (grc), and Latin (lat)—with a primary focus on Sanskrit. These languages represent a unique category of low-resource languages: despite scarce data for downstream NLU tasks, they have rich ancient literature available in digitized formats (Goyal et al., 2012; Berti, 2019) and exert significant influence on the vocabulary and narrative structures of better-resourced languages (e.g., Latin contributes approximately 28% of English vocabulary (Finkenstaedt and Wolff, 1973)). Moreover, the high inflection these languages present a challenge for evaluating LLM generalization.
To investigate these issues, we conduct two sets of zero-shot experiments using gpt-4o (OpenAI, 2024, 2023), llama-3.1-405b-instruct (Dubey et al., 2024), and their smaller variants. First, we assess performance on two NLU tasks with available datasets for all three languages, namely, named entity recognition (NER) and machine translation to English (MT). We observe that larger models perform comparably or even better than previously reported fine-tuned models on out-of-domain data. Second, we focus on Sanskrit by introducing a factoid closed-book QA dataset and show that question-answering performance improves with retrieval augmented generation (RAG) (Lewis et al., 2020) when the model is provided with relevant texts. Given the recent nature of these datasets relative to the models’ knowledge cut-off dates—and the likelihood that instruction-tuning on these languages is limited—the robust performance observed can be attributed to cross-lingual generalization. In both experimental setups, we find that smaller models struggle, particularly with niche entity types in NER, and in effectively leveraging contextual information via RAG suggesting model scale as an important factor.
2 Datasets and Methods
Task Lang. Source Size
NER san Terdalkar (2023) 139
lat Erdmann et al. (2019) 3410
grc Myerston (2025) 4957
MT-en san Maheshwari et al. (2024) 6464
lat Rosenthal (2023) 1014
grc Palladino et al. (2023) 274
QA san Ours 1501
Table 1: An overview of NLU datasets used in this study for Sanskrit (san), Latin (lat) and Ancient Greek (grc). QA dataset for san is contributed in this work. Size represents the number of sentences of test sets (wherever train-test splits exist).
The datasets used in our experiments are summarized in Table 1. Our aim is to evaluate zero-shot capabilities where evaluation is done directly on test data without fine-tuning on the training data. Thus, we only consider the test sets of these datasets. Notably, the Sanskrit NER dataset (san) is the smallest, comprising 139 sentences (1558 tokens) (Terdalkar, 2023). In addition to these publicly available datasets, we contribute a new factoid closed-domain QA dataset in Sanskrit, described in detail in Section 2.1.
We evaluate the zero-shot capabilities of both large and small variants of our models: proprietary (gpt-4o and gpt-4o-mini (OpenAI, 2024)) and open-source (llama-3.1-405b-instruct and llama-3.1-8b-instruct (Dubey et al., 2024)). According to official documentation, these models have knowledge cut-off dates at the end of 2023. Given that none of the documentation indicates explicit instruction tuning on Sanskrit, Ancient Greek, or Latin, any observed performance in these languages is attributable to cross-lingual generalization. Previous zero-shot applications of LLMs to classical languages have been limited to Latin machine translation and summarization (Volk et al., 2024). All prompts used for these tasks are provided in Appendix C. The prompts are in English but also tested with manually translated san prompts for san. For the NER task, the prompt includes a single example to illustrate the output format, though it is effectively zero-shot for most entity types, as the example contains at most two.
2.1 Sanskrit QA
To further evaluate comprehension, we focus on question-answering (QA) in Sanskrit—a domain with very limited datasets. The only existing dataset by Terdalkar and Bhattacharya (2019) comprises 80 kinship-related questions. To address this gap, we created a new dataset containing 1501 factoid QA pairs covering distinct domains in Sanskrit: epics and healthcare. Specifically, we selected two key texts: (1) Rāmāyaṇa, an ancient Indian epic, and (2) Bhāvaprakāśanighaṇṭu, a foundational text on Āyurveda. Details of the dataset are provided in Appendix A.
For QA evaluation, we employ a closed-book setting using prompts detailed in Appendix C.3. Additionally, to simulate extractive QA, we implement Retrieval Augmented Generation (RAG) by retrieving the top-$k$ relevant passages (with $k$ tuned to 4) from the original Sanskrit texts using BM25 retrieval (Sparck Jones, 1972; Robertson et al., 2009). We found BM25 based RAG (RAG-BM25) to outperform RAG based on other possible retrievers that employ word-embeddings namely, FastText (Bojanowski et al., 2017) and GloVe (Pennington et al., 2014) (reported in Appendix B). To determine if the presence of the answer in the context benefits the LLMs, we manually annotated the relevance of the retrieved passages. Since BM25 computes similarities between queries and texts preferably based on lemmas rather than inflected forms, a lemmatizer is needed in the pre-processing step. For this purpose, we trained a Seq2Seq transformer-based Sanskrit lemmatizer based on the DCS corpus (Hellwig, 2010-2024), achieving a mean F1-score of 0.94 on a held-out test set. Implementation details are given in Appendix D. The code and data are available at https://github.com/mahesh-ak/SktQA.
3 Results
Figure 1: Zero-shot evaluation of LLMs on three NLU tasks for classical languages (language codes in ISO 639-2). Prompts are in English except that <san> represents prompt in san. Larger LLMs are represented in red and orange, while smaller LLMs in blue and purple. First row displays the performances on NER (a) and MT (to en) (b) for all three languages. Second row displays the performances on QA for Sanskrit alone. Out of 1501 QA pairs whose chunks are considered while plotting (c), there are 607 QA pairs with answer present in at least one of the $k=4$ contexts extracted by BM25 and 894 QA pairs with answer not inferable from contexts, which are respectively considered in (d) and (e).
Figure 1 presents our zero-shot evaluation results, demonstrating that larger LLMs exhibit robust cross-lingual generalization across four NLU tasks—named entity recognition (NER), machine translation (MT), closed-book QA, and extractive QA (simulated via RAG-BM25)—in three classical languages (with QA evaluated solely on Sanskrit). To assess variability, each test set is segmented into 10 chunks during evaluation resulting in a box-plot (Fig. 1). Larger models perform comparable or better than previous fine-tuned models on out-of-domain data as reported in Appendix E. Notably, when answer-bearing contexts are provided (Fig. 1d) versus when they are absent (Fig. 1e), the models show significant performance gains, suggesting comprehension abilities in Sanskrit, a language characterized by high inflection. This behavior is however, not exhibited by smaller models when prompted in Sanskrit.
3.1 Sanskrit vs English Prompts
In our evaluations for san, we prompt in both English and Sanskrit, as mentioned earlier. Figure 1 shows that, for most models, prompting in English significantly outperforms prompting in Sanskrit—except for gpt-4o, where performance is largely similar. This provides partial evidence that these models are not instruction tuned on Sanskrit, aligning with trends observed by Muennighoff et al. (2023). In the case of gpt-4o, the comparable performance may indicate that the model has reached a saturation point to its best possible zero-shot performance as it is unlikely to have been trained on substantially more instruction tuning data than gpt-4o-mini.
Figure 2: Confusion matrices from the NER task in san (a-d), lat (e-h) and grc (i-l), normalized across rows.
3.2 Misclassified Entities in NER
Figure 2 displays confusion matrices for the NER task. Across all three languages, the smaller models exhibit more confusion among semantically related classes, while the larger models show fewer off-diagonal errors. In san, mythological entities like DEVA, ASURA, and RAKSHASA or semantically closed entities like KINGDOM vs CITY (eg., kośala vs ayodhyā) or FOREST (eg., daṇḍaka) vs GARDEN (eg., nandana, a heavenly garden)often get misclassified with each other in the smaller models. For lat, entity type GRP proves troublesome for the smaller models, suggesting struggles in separating individual entities from collective ones. In grc, categories such as LOC and ORG show higher confusion counts akin to GRP in lat while confusion between GOD and PERSON seems similar to what we have just discussed in Sanskrit. In contrast, much clearer boundaries emerge in the larger models’ confusion matrices. In many of these cases, the domain or style of the texts—especially if they involve mythological or archaic terms typical of classical texts— can be understood to influence performance, as models with less exposure to specialized terminology may conflate related entity types.
3.3 Inflection in Answers of san QA
LLM Closed-book + RAG-BM25
Inflected Lemmatized Inflected Lemmatized
gpt-4o 0.36 0.37 0.46 0.48
llama-3.1-405b-instruct 0.41 0.40 0.42 0.44
gpt-4o-mini 0.18 0.20 0.25 0.28
llama-3.1-8b-instruct 0.13 0.15 0.09 0.10
Table 2: Comparison of EM scores in san QA (English prompts) when predicted and gold answers are considered with inflection or lemmatized.
In the Sanskrit QA task, models are expected to generate single-word answers with correct inflection. For computing exact match (EM) scores, we manually identified all acceptable answers, excluding those with incorrect inflection (e.g., wrong case or gender endings). To quantify inflection errors, we also computed EM scores on lemmatized versions of both the gold-standard and predicted answers, as shown in Table 2. Most models exhibit only a slight increase in EM scores on lemmatized answers, suggesting that inflection errors are relatively minor—a finding corroborated by manual inspection. Future work could extend this analysis to investigate inflection accuracy in full sentence generation within broader Natural Language Generation scenarios.
3.4 Knowledge-Graph QA
Additionally, we explored the use of incomplete knowledge graphs (KGs) for Sanskrit QA. We evaluated a KG derived from the Bhāvaprakāśanighaṇṭu text (Terdalkar et al., 2023) and constructed a small KG for Rāmāyaṇa (details in Appendix F). Using the Think-On-Graph (ToG) paradigm (Sun et al., 2024), which iteratively explores KG paths for answer retrieval in a zero-shot, training-free manner (Xu et al., 2024), we observed that gpt-4o could effectively execute this method. Although it occasionally extracted correct answers, its performance did not significantly exceed that of the closed-book setting, possibly due to the incompleteness of the KGs. Future work may focus on developing more comprehensive KGs to enhance this retrieval method.
4 Conclusions
In summary, our zero-shot evaluations demonstrate that larger language models exhibit robust cross-lingual generalization across diverse natural language understanding tasks in classical languages—including NER, machine translation, and QA. Notably, the significant performance gains achieved when answer-bearing contexts are provided—particularly in Sanskrit QA—suggest comprehension abilities in highly inflected languages. Moreover, our contribution of a novel Sanskrit QA dataset provides a valuable resource for evaluating and benchmarking LLM performance on classical language tasks. Importantly, these models have not been explicitly instruction tuned on Sanskrit, Latin, or Ancient Greek—evidenced by the superior performance achieved when using English prompts for Sanskrit—which indicates that their zero-shot performance is attributable solely to cross-lingual generalization. Future work should focus on expanding dataset coverage and exploring additional classical languages and tasks, further advancing our understanding of cross-lingual generalization in LLMs and its applications in digital humanities and multilingual NLP research.
Limitations
While our study demonstrates robust cross-lingual generalization in large language models for classical languages, several limitations warrant discussion. First, our newly contributed Sanskrit QA dataset, although valuable, is limited in size. Our evaluation relies exclusively on zero-shot performance, as the models have not been explicitly instruction tuned on these languages; this design choice may obscure potential benefits achievable through targeted fine-tuning. Furthermore, the effectiveness of our BM25-based retrieval approach depends heavily on preprocessing steps such as lemmatization, which might not optimally address all linguistic variations in highly inflected languages. Finally, our comparisons are based on a limited set of proprietary and open-source models, and future work should extend this analysis to a broader range of models and tasks to fully understand the nuances of cross-lingual generalization in classical languages.
Ethics Statement
Classical Sanskrit epics hold deep cultural and religious significance in Indian traditions, and similarly, Āyurveda represents a sensitive area within healthcare. We acknowledge that any research involving these subjects must be conducted with particular care. It is essential to note that, as with conventional treatment, Āyurvedic practices require professional consultation and should not be substituted by automated responses. Although our experiments indicate that paradigms like RAG produce more grounded and, hence, potentially safer outputs, there is no assurance that the responses from current LLMs in these domains meet clinical or religious safety standards. Accordingly, the authors do not endorse using the datasets beyond the scope of linguistic research. These datasets will be released under an open-source, non-commercial license, and all annotators have been compensated at fair, standard rates.
Appendix A Question Answering Dataset
In this appendix, we describe the creation of Sanskrit QA dataset.
We referred to two books that contain multiple-choice questions (MCQs) with answers: one comprising 1000 MCQs on Rāmāyaṇa Singh (2009), and another featuring a collection of 2600 questions from three prominent texts of Āyurveda Phull and Phull (2017). The questions and options in these books are in Hindi language.
We carefully selected a relevant subset of questions from these books, including all 1000 questions from Rāmāyaṇa dataset and $431$ from that of Āyurveda. These questions are then translated into Sanskrit with the help of experts in the language who are also familiar with the original Sanskrit texts. Further, we consulted with a specialist in Āyurveda to review and discard incorrect question-answer pairs, as well as to generate 70 new questions based on Bhāvaprakāśanighaṇṭu. Ultimately, the question-answering dataset consists of $1501$ questions.
The answers typically agree in grammatical case with the corresponding interrogative of the question. The following is a question-answer pair as an illustration111gen - genitive, loc - locative:
Q: śītala-jalasya pānaṃ kasmin roge niṣiddham asti? A: gala-grahe
Q: cold-water.gen drinking what.loc disease.loc forbidden is A: pharyngitis.loc
Question: During which condition is the drinking of cold water forbidden? Answer: During pharyngitis.
Most questions in the datasets have a single-word answer except a few including those in the Rāmāyaṇa that fall under the category ‘Origins’ (Table 4). An example question-answer pair under this category that demands reasoning in the answer:
Q: rājā-sagareṇa sagaraḥ iti nāma kutaḥ prāptam?
“How did King Sagara obtain such a name?”
A: saha tena gareṇaiva jātaḥ sa sagaro ’bhavat
“He was indeed born along with (sa-) the poison (gara), thus he became Sagara.”
For such questions (only about 50), the answers can be paraphrased variously, thereby requiring manual evaluation.
The broad semantic and domain-specific categories of the questions are detailed in Tables 4 and 4.
Category Description #Q
Names Names of various characters 97
Actions Who performed certain actions? 47
Origins Origin of various names 49
Numeric Questions with numerical answers 79
Quotes Who said to whom? 31
Boons and Curses Who endowed boons / curses on whom 31
Weapons Questions related to various types of weapons 59
Locations Locations of important events or characters 71
Kinship Questions pertaining to human kinship relationships 133
Slay Who slayed whom 49
Kingdoms Which king ruled which kingdom 27
Incarnations Who were incarnations of which deities 27
MCQ Multiple choice questions 140
Miscellaneous Other questions 196
Table 3: Question Categories for Rāmāyaṇa QA Dataset
Category Description #Q
Synonym Synonyms of substances 174
Type Variants or types of substances 30
Property Properties of substances 20
Comparison Comparison between properties of various substances or their variants 24
Consumption Related to consumption of medicine including suitability, method, accompaniments etc. 23
Count Counting types or properties of substances 59
Quantity Quantity of substances in various procedures or methods 21
Time-Location Time or location in the context of substances or methods 17
Effect Effect of substances 15
Treatment Diseases and treatments 23
Method Methods of preparation of substances 21
Meta Related to the verbatim source text, the structure of the text and external references 38
Multi-Concept About more than one aforementioned concepts 11
Miscellaneous Miscellaneous concepts 24
Table 4: Question Categories for Āyurveda QA Dataset
Appendix B Retrieval Augmented Generation
In the RAG paradigm, the LLM is provided with additional context that consists of top-$k$ passages retrieved from the original texts. The texts of Rāmāyaṇa and Bhāvaprakāśanighaṇṭu are obtained from GRETIL [^2] https://gretil.sub.uni-goettingen.de/ and Sanskrit Wikisource [^3] https://sa.wikisource.org/wiki/ respectively. The texts are pre-processed following standard procedures (Manning, 2008), namely, dividing the texts into chunks, followed by lemmatization, and then building a document store. Lemmatization would not have been necessary if retrieval frameworks such as Dense Passage Retrieval (Karpukhin et al., 2020) or a vector space retrieval framework with SentenceBERT embeddings (Reimers and Gurevych, 2019) could be used. However, due to insufficient data in Sanskrit, such models cannot be trained now. Hence, we used BM25 retrieval and vector space retrieval with averaged FastText (AvgFT) (Bojanowski et al., 2017) and GloVe (Pennington et al., 2014) (AvgGV) embeddings, which are employed on lemmatized documents and queries. To achieve this, a lemmatizer for Sanskrit was built as described below.
Sanskrit Lemmatizer
Seq2Seq transformer-based Sanskrit lemmatizer was trained from the words and their respective lemmas present in the DCS corpus (Hellwig, 2010-2024) [^4] http://www.sanskrit-linguistics.org/dcs/. During lemmatization, if a word in an input sentence is a compound word or involves Sandhi, the lemmatizer is expected to break the word into sub-words and generate their respective lemmas in the output. For example, if the input sentence is ‘haridrāmalakaṃ gṛhṇāti’, the corresponding lemmatized output should be ‘haridrā āmalaka gṛh’. Our lemmatizer achieves a mean F1-score of 0.94 across the sentences from the held-out test set (Appx. D) calculated according to Melamed et al. (2003), however with a significant standard deviation of 0.11. While the accuracy is high, future attempts for improvements should focus on minimizing the variance, which is rarely ever reported although important.
The information retrieval pipelines thus formulated can be considered novel concerning Classical Sanskrit. A known earlier attempt towards building retrieval systems in Sanskrit (Sahu and Pal, 2023) focused on news corpora with much terminology consisting of borrowings from Hindi and even English. As a result, the lemmatizer trained on Classical Sanskrit and thereby, our entire retrieval pipeline may not be appropriate on such corpora and hence are not comparable.
We evaluated RAG with different $k$ using GPT-4o with prompts in Sanskrit. The EM scores are plotted in Fig. 3. It can be seen that BM25 outperforms AvgFT and AvgGV, it may also be inferred from both plots that $k=4$ would be optimal. At this value of $k$, the performance of RAG in the case of Bhāvaprakāśanighaṇṭu already starts to deteriorate, although not significantly. Keeping both plots in mind, we have fixed the value of $k$ at 4.
Figure 3: Effect of $k$ on RAG, denoting the number of top best matching text chunks retrieved, on the performances of GPT-4o with retrievers based on BM25, averaged FastText (AvgFT) and GloVe (AvgGV) embeddings respectively of datasets Rāmāyaṇa (left) and Bhāvaprakāśanighaṇṭu (right).
The prompts for RAG are detailed in Appx. C.4.
Appendix C Prompts
The Sanskrit prompts are in Devanāgari script. In this appendix, we provide the prompts (transliterated in IAST scheme [^5] https://en.wikipedia.org/wiki/International_Alphabet_of_Sanskrit_Transliteration).
C.1 Prompts for Named Entity Recognition
Prompt in English
Recognize the named entities from the following sentence in {LANGUAGE}. The valid tags are {ENTITY TYPES}. Do not provide explanation and do not list out entries of ‘O’.
Example:{EXAMPLE}
(Note: EXAMPLE in Sanskrit is:
daśarathaḥ ayodhyāyāḥ rājā āsīt
{‘B-HUMAN’: [‘daśarathaḥ’], ‘B-CITY’: [‘ayodhyāyāḥ’]}
An example is provided to make the output structure clear.)
Prompt in Sanskrit
adho datta vākye named entities abhijānīhi (NER kuru). tadapi vivṛtam mā kuru, kevalam pṛṣṭa viṣayasya uttaram dehi.
entities eteṣu vargeṣu vartante - {ENTITY TYPES}
udāharaṇāya - {EXAMPLE}
C.2 Prompts for Machine Translation
Prompt in English
Translate the following sentence in {LANGUAGE} into English. Do not give any explanations.
Prompt in Sanskrit
adho datta-saṃskṛta-vākyam āṅgle anuvādaya, tad api vivṛtam mā kuru -
(Sanskrit QA Prompts)
In the following prompts, TOPIC is either ‘Rāmāyaṇa’ or ‘Āyurveda’.
C.3 Prompts for Closed-book QA
Prompt in English
Answer the question related to {TOPIC} in the Sanskrit only. Give a single word answer if reasoning is not demanded in the answer. With regards to how-questions, answer in a short phrase, there is no single word restriction.
{QUESTION} {CHOICES}
Prompt in Sanskrit
tvayā saṃskṛta-bhāṣāyām eva vaktavyam. na tu anyāsu bhāṣāsu. adhaḥ {TOPIC}-sambandhe pṛṣṭa-praśnasya pratyuttaraṃ dehi. tadapi ekenaiva padena yadi uttare kāraṇam nāpekṣitam. katham kimartham ityādiṣu ekena laghu vākyena uttaraṃ dehi atra tu eka-pada-niyamaḥ nāsti.
{QUESTION} {CHOICES}
C.4 Prompts for RAG-QA
Prompt in English
Answer the following question related to {TOPIC} in Sanskrit only. Give a single word answer if reasoning is not demanded in the answer. With regards to how-questions, answer in a short phrase. Also take help from the contexts provided. The contexts may not always be relevant."
contexts: {CONTEXTS}
question:{QUESTION} {CHOICES}
Prompt in Sanskrit
tvayā saṃskṛta-bhāṣāyām eva vaktavyam. na tu anyāsu bhāṣāsu. adhaḥ {TOPIC}-sambandhe pṛṣṭa-praśnasya pratyuttaraṃ dehi. tadapi ekenaiva padena, yāvad laghu śakyaṃ tāvad, taṃ punaḥ vivṛtam mā kuru. api ca yathā’vaśyam adhaḥ datta-sandarbhebhyaḥ ekatamāt sahāyyaṃ gṛhāṇa. tattu sarvadā sādhu iti nā’sti pratītiḥ.
sandarbhāḥ: {CONTEXTS}
praśnaḥ: {QUESTION} {CHOICES}
Appendix D Implementation
This appendix outlines the implementation details. All LLMs are operated through API calls using LangChain [^6] https://www.langchain.com/. In case of Llama-3.1, we used API provided by Fireworks AI [^7] https://fireworks.ai/.
The lemmatizer was implemented using HuggingFace transformers (Wolf et al., 2020) upon base model T5 (Raffel et al., 2020) initiated with the model configuration of 4 layers per each encoder and decoder, 4 attention-heads, embedding of size 256, and hidden size of 1024, totaling about 100M parameters. The tokenizer trained by Akavarapu and Bhattacharya (2023) was used [^8] https://huggingface.co/mahesh27/vedicberta-base. The lemmatizer was trained for 15 epochs on DCS (Hellwig, 2010-2024) data with batch size of 32, that took about 15 hours on NVIDIA RTX 2080 with 11GB graphics memory. There are total 1.04M sentences in the data, that are randomly divided into proportions $0.675:0.075:0.15$ respectively for training, validation and testing. FastText and GloVe embeddings are trained on lemmas obtained from DCS (Hellwig, 2010-2024) with embedding size 100.
Appendix E Supplementary Results
Model BLEU
Google Trans (Maheshwari et al., 2024) 13.9
IndicTrans (Maheshwari et al., 2024) 13.1
gpt-4o 16.5
llama-3.1-405b-instruct 17.1
MT (san-eng)
Model Macro F1 (BI)
LatinBERT1 (Beersmans et al., 2023) 0.54
LatinBERT2 (Beersmans et al., 2023) 0.50
gpt-4o 0.51
llama-3.1-405b-instruct 0.49
NER (lat)
Table 5: Comparison of out of domain performances of LLMs against previously reported fine-tuned models.
In Table 5, we compare the out-of-domain performance of our evaluated models against previously reported fine-tuned models. For MT (san-eng), open-source model llama-3.1-405b-instruct achieves a BLEU score of 17.1, outperforming both Google Trans and IndicTrans, while for NER (lat), the performance of LLMs is roughly comparable to that of fine-tuned LatinBERT variants. Although fine-tuned models yield superior results on in-domain data, our findings indicate that multilingual LLMs are competitive—or even superior—in their zero-shot generalization.
Appendix F LLMs with Knowledge Graphs
Arriving at an answer by an LLM integrated with a knowledge graph (KG) through Think-on-Graph (ToG) (Sun et al., 2024) paradigm involves several prompting steps for each hop from starting entity nodes. Firstly, the LLM lists entities from the input questions further lemmatized by our lemmatizer previously described. The relationships from and to these entities are then extracted by traversing the KG. The LLM then lists relationships with relevance scores, which are further used to prune the relationships, retaining only the best three. Unexplored entities connected by these relationships are then known from the KG, which are similarly pruned to retain the three most relevant ones. The LLM then reasons whether these extracted paths suffice to answer the given question. If no, the cycle is repeated, i.e., it traverses a hop further up to a depth $d$. Otherwise, the LLM answers using the context from the extracted paths.
The prompts for each step and an outline pseudo-code can be found respectively in Appx. F.2, Alg. 1. Technical terminology such as ‘entity’, ‘knowledge graph’, and so forth are mostly retained in English in these prompts resulting in minimal and unavoidable code-mixing. Further, the output of these prompts is often a list of elements and, hence, has to abide by a structured format.
F.1 Knowledge Graphs
A knowledge graph (KG) was constructed for Rāmāyaṇa using two key references, Ray (1984) and Rai (1965). The graph was annotated with the help of two experts proficient in both Sanskrit and Rāmāyaṇa. For annotation, we used a custom deployment of Sangrahaka Terdalkar and Bhattacharya (2021). The resulting knowledge graph contains 867 nodes and 944  relations, encompassing entities like characters of the story including humans and divine beings, places (cities, rivers, kingdoms), and animals, and relationships such as kinship, actions, locations, and others, highlighting associations between the characters, natural features, and other elements from the text.
Additionally, a work-in-progress knowledge graph for Bhāvaprakāśanighaṇṭu obtained from the authors of Terdalkar et al. (2023) was referenced. The KG currently includes 4685 nodes and 10596 relations from 12 out of 23 chapters covering substances such as grains, vegetables, meats, metals, poisons, dairy products, prepared substances and other miscellaneous medicinal substances.
The knowledge graphs were loaded and accessed through Neo4j [^9] https://neo4j.com/. Python packaage, indic-transliteration [^10] https://github.com/indic-transliteration/indic_transliteration_py is used to move among transliteration schemes of Sanskrit. The pseudo-code for our implementation of ToG (Sun et al., 2024) is given in Algorithm 1. The sample limit $S$ is set to $15$, depth limit $D$ to $1$ and width limit $W$ to 3.
Algorithm 1 Outline of LLM-KG i.e., ToG (Sun et al., 2024)
Input: $x$
LLM prompt-chains: ExtractEntities, RelationPrune, EntityExtractPrune, Reason, Answer
Interface to KG: FetchRelations, FetchEntities; Depth limit: $D$; Sample limit for KG: $N$; Width limit for LLM: $W$
Current Entities $E\leftarrow$ ExtractEntities($x$)
Current depth $d\leftarrow 0$
Stored Paths $P\leftarrow$ []
while $d<D$ do
     $R\leftarrow$ FetchRelations($E$, $N$)
     $R\leftarrow$ RelationPrune($R$, $W$)
     $E,P\leftarrow$ FetchEntities($E$, $R$, $P$, $N$)
     $E,P\leftarrow$ EntityExtractPrune($E$, $R$, $P$, $W$)
     if Reason($x$, $E$, $P$then
         Answer($x$, $E$, $P$)
         break
     end if
     $d\leftarrow d+1$
end while
if $d=D$ then Answer($x$, $E$, $P$)
end if
F.2 LLM-KG Prompts
ExtractEntities
system tvam knowledge-graph-taḥ uttarāṇi niṣkarṣyituṃ praśnāt entities vindasi ca tāni saha relevance-score (0-1 madhye) samarpayasi.
output udāharaṇam (‘rāmaḥ’, 0.8), (‘sītā’, 0.7). tato vivṛtaṃ mā kuru.
human praśnaḥ: {QUESTION} {CHOICES}
RelationPrune
system tvam datta-praśnasya uttarāṇi knowledge-graph-taḥ niṣkarṣituṃ knowledge-graph-taḥ idānīṃ paryantaṃ niṣkarṣita-sambandhebhyaḥ avaśyāni saha relevance-score (0-1 madhye) samarpayasi.
output udāharaṇam (‘IS_FATHER_OF’, 0.8), (‘IS_CROSSED_BY’, 0.7), $\ldots$ . tato vivṛtaṃ mā kuru.
human praśnaḥ: {QUESTION} {CHOICES}
niṣkarṣitāni sambandhāni: {RELATIONS}
EntityExtractPrune
system tvam datta-praśnasya uttarāṇi knowledge-graph-taḥ niṣkarṣituṃ knowledge-graph-taḥ idānīṃ paryantaṃ niṣkarṣita-sambandhebhyaḥ avaśyāni nodes (lemmas) saha relevance-score (0-1 madhye) samarpayasi.
output udāharaṇam (‘rāmaḥ’, 0.8), (‘sītā’, 0.7). tato vivṛtaṃ mā kuru.
human praśnaḥ: {QUESTION} {CHOICES}
niṣkarṣitāni sambandhāni: {RELATIONS, ENTITIES}
Reason
system tvam datta-praśnasya uttarāṇi knowledge-graph-taḥ niṣkarṣituṃ knowledge-graph-taḥ idānīṃ paryantaṃ niṣkarṣitaṃ yat-kiñcid praśnasya uttaraṃ dātuṃ alam (1) vā nālam (0) iti vaktavyam.
output 1 athavā 0. na anyat vadasi
human praśnaḥ: {QUESTION} {CHOICES}
niṣkarṣitam: {PATHS}
Answer
system adhaḥ {TOPIC}-sambandhe pṛṣṭa-praśnasya pratyuttaraṃ dehi. tadapi praśnocitavibhaktau bhavet na tu prātipadika rūpe. tadapi ekenaiva padena yadi uttare kāraṇam nāpekṣitam. katham kimartham ityādiṣu ekena laghu vākyena uttaraṃ dehi atra tu eka-pada-niyamaḥ nāsti.
api ca yathā’vaśyam adhaḥ dattaiḥ knowledge-graph-taḥ niṣkarṣita-viṣayaiḥ sahạyyaṃ gṛhāṇa. tattu sarvadā sādhu iti nā’sti pratītiḥ. uttaram yāvad laghu śakyam tāvat laghu bhavet.
human praśnaḥ: {QUESTION} {CHOICES}
niṣkarṣitam: {PATHS}
uttaram:

### v2

@@
 Previous work has demonstrated generalization to several low-resource languages via few-shot in-context learning (Cahyawijaya et al., 2024).
-In this study, we focus on zero-shot generalization to natural language understanding (NLU) tasks for three Indo-European classical languages—Sanskrit (san), Ancient Greek (grc), and Latin (lat)—with a primary focus on Sanskrit.
-These languages represent a unique category of low-resource languages: despite scarce data for downstream NLU tasks, they have rich ancient literature available in digitized formats (Goyal et al., 2012; Berti, 2019) and exert significant influence on the vocabulary and narrative structures of better-resourced languages (e.g., Latin contributes approximately 28% of English vocabulary (Finkenstaedt and Wolff, 1973)).
-Moreover, the high inflection these languages present a challenge for evaluating LLM generalization.
-To investigate these issues, we conduct two sets of zero-shot experiments using gpt-4o (OpenAI, 2024, 2023), llama-3.1-405b-instruct (Dubey et al., 2024), and their smaller variants.
+In this study, we focus on zero-shot generalization to natural language understanding (NLU) tasks for three classical languages—Sanskrit (san), Ancient Greek (grc), and Latin (lat)—with a primary focus on Sanskrit.
+These languages represent a unique category of low-resource languages – despite scarce data for downstream NLU tasks, they have rich ancient literature available in digitized formats (Goyal et al., 2012; Berti, 2019), and they exert significant influence on the vocabulary and narrative structures of better-resourced languages (e.g., Latin contributes approximately 28% of English vocabulary (Finkenstaedt and Wolff, 1973)).
+Moreover, the high inflection of these languages presents a challenge.
+Figure 1: The three NLU tasks evaluated on the classical languages: Named-Entity Recognition (top-left), Machine Translation (bottom-left) and Question-Answering (right).
+To investigate these issues, we conduct two sets of zero-shot experiments using gpt-4o (OpenAI, 2024; OpenAI et al., 2023), llama-3.1-405b-instruct (Dubey et al., 2024), and their smaller variants.
 First, we assess performance on two NLU tasks with available datasets for all three languages, namely, named entity recognition (NER) and machine translation to English (MT).
@@
 Second, we focus on Sanskrit by introducing a factoid closed-book QA dataset and show that question-answering performance improves with retrieval augmented generation (RAG) (Lewis et al., 2020) when the model is provided with relevant texts.
-Given the recent nature of these datasets relative to the models’ knowledge cut-off dates—and the likelihood that instruction-tuning on these languages is limited—the robust performance observed can be attributed to cross-lingual generalization.
-In both experimental setups, we find that smaller models struggle, particularly with niche entity types in NER, and in effectively leveraging contextual information via RAG suggesting model scale as an important factor.
+The tasks are illustrated in Figure 1.
+Given the recent nature of these datasets relative to the models’ knowledge cut-off dates, and the likelihood that instruction-tuning on these languages is limited, the robust performance observed can be attributed to cross-lingual generalization.
+We refer to our prompting strategy as zero-shot, as it includes no task-specific examples, and it is unlikely that such examples in these languages were present in the models’ training or instruction-tuning data.
+In both experimental setups, we find that smaller models struggle, particularly with niche entity types in NER, and in effectively leveraging contextual information via RAG, thereby suggesting model scale as an important factor.
 2 Datasets and Methods
-Task Lang. Source Size
+Task Language Source Size
 NER san Terdalkar (2023) 139
@@
 According to official documentation, these models have knowledge cut-off dates at the end of 2023.
-Given that none of the documentation indicates explicit instruction tuning on Sanskrit, Ancient Greek, or Latin, any observed performance in these languages is attributable to cross-lingual generalization.
-Previous zero-shot applications of LLMs to classical languages have been limited to Latin machine translation and summarization (Volk et al., 2024).
-All prompts used for these tasks are provided in Appendix C.
-The prompts are in English but also tested with manually translated san prompts for san.
-For the NER task, the prompt includes a single example to illustrate the output format, though it is effectively zero-shot for most entity types, as the example contains at most two.
+Many datasets considered in this work (Table 1) are released beyond these dates, in other words, they are unlikely to be seen in the pre-training data of these models.
+Given that none of the documentation indicates explicit instruction tuning on Sanskrit, Ancient Greek, or Latin, any observed performance in these languages is likely attributable to cross-lingual generalization.
+Previous zero-shot applications of LLMs to classical languages have been limited to Latin machine translation and summarization (Volk et al., 2024), although several works have been dedicated to building language models for these languages (Riemenschneider and Frank, 2023; Nehrdich et al., 2024), however, with fine-tuning restricted to morphological parsing-related tasks like dependency parsing (Nehrdich and Hellwig, 2022; Hellwig et al., 2023; Sandhan et al., 2023).
+All prompts used for these tasks are provided in Appendix A. The prompts are tested in both English and the respective languages.
 2.1 Sanskrit QA
-To further evaluate comprehension, we focus on question-answering (QA) in Sanskrit—a domain with very limited datasets.
+To further evaluate comprehension, we focus on question-answering (QA) in Sanskrit – a domain with very limited datasets.
 The only existing dataset by Terdalkar and Bhattacharya (2019) comprises 80 kinship-related questions.
@@
 Specifically, we selected two key texts: (1) Rāmāyaṇa, an ancient Indian epic, and (2) Bhāvaprakāśanighaṇṭu, a foundational text on Āyurveda.
-Details of the dataset are provided in Appendix A.
-For QA evaluation, we employ a closed-book setting using prompts detailed in Appendix C.3.
-Additionally, to simulate extractive QA, we implement Retrieval Augmented Generation (RAG) by retrieving the top-$k$ relevant passages (with $k$ tuned to 4) from the original Sanskrit texts using BM25 retrieval (Sparck Jones, 1972; Robertson et al., 2009).
-We found BM25 based RAG (RAG-BM25) to outperform RAG based on other possible retrievers that employ word-embeddings namely, FastText (Bojanowski et al., 2017) and GloVe (Pennington et al., 2014) (reported in Appendix B).
-To determine if the presence of the answer in the context benefits the LLMs, we manually annotated the relevance of the retrieved passages.
-Since BM25 computes similarities between queries and texts preferably based on lemmas rather than inflected forms, a lemmatizer is needed in the pre-processing step.
-For this purpose, we trained a Seq2Seq transformer-based Sanskrit lemmatizer based on the DCS corpus (Hellwig, 2010-2024), achieving a mean F1-score of 0.94 on a held-out test set.
-Implementation details are given in Appendix D.
-The code and data are available at https://github.com/mahesh-ak/SktQA.
+Details of the dataset are provided in Appendix B.
+Figure 2: Effect of $k$ on RAG, denoting the number of top best matching text chunks retrieved, on the performances of GPT-4o with retrievers based on BM25, averaged FastText (AvgFT) and GloVe (AvgGV) embeddings respectively of datasets Rāmāyaṇa (left) and Bhāvaprakāśanighaṇṭu (right).
+For QA evaluation, we employ a closed-book setting using prompts detailed in Appendix A.3.
+To emulate extractive QA, we implement a Retrieval-Augmented Generation (RAG) approach by retrieving the top-$k$ relevant passages ($k$ tuned to 4) from the original Sanskrit texts using BM25 (Sparck Jones, 1972; Robertson et al., 2009).
+We also compare BM25 with embedding-based retrievers—FastText (Bojanowski et al., 2017) and GloVe (Pennington et al., 2014)—and vary $k$ to assess performance using gpt-4o with Sanskrit prompts.
+As shown in Fig.
+2, BM25 consistently outperforms the embedding-based methods, and $k=4$ emerges as an optimal choice across metrics.
+To examine whether the inclusion of answer-bearing contexts benefits model performance, we manually annotated the relevance of retrieved passages.
+Since BM25 relies on lexical similarity, typically favoring lemmas over inflected forms, we introduce a lemmatization step using a transformer-based Seq2Seq Sanskrit lemmatizer trained on the DCS corpus (Hellwig, 2010-2024), achieving a mean F1 score of 0.94 on a held-out test set.
+Further details on RAG and lemmatization are provided in Appendix C, and implementation details in Appendix D.
+Code and data are available at https://github.com/mahesh-ak/SktQA.
 3 Results
 Figure 3: Zero-shot evaluation of LLMs on three NLU tasks for classical languages (language codes in ISO 639-2).
-Prompts are in English except that <san> represents prompt in san.
+Prompts when in English are denoted by <en>, otherwise are in respective languages.
 Larger LLMs are represented in red and orange, while smaller LLMs in blue and purple.
@@
 Second row displays the performances on QA for Sanskrit alone.
-Out of 1501 QA pairs whose chunks are considered while plotting (c), there are 607 QA pairs with answer present in at least one of the $k=4$ contexts extracted by BM25 and 894 QA pairs with answer not inferable from contexts, which are respectively considered in (d) and (e).
+Out of 1501 QA pairs considered (c), 607 QA pairs are with answer present in at least one of the $k=4$ contexts extracted by BM25 and 894 QA pairs with answer not inferable from contexts, which are respectively considered in (d) and (e).
 Figure 3 presents our zero-shot evaluation results, demonstrating that larger LLMs exhibit robust cross-lingual generalization across four NLU tasks—named entity recognition (NER), machine translation (MT), closed-book QA, and extractive QA (simulated via RAG-BM25)—in three classical languages (with QA evaluated solely on Sanskrit).
-To assess variability, each test set is segmented into 10 chunks during evaluation resulting in a box-plot (Fig.
-1).
-Larger models perform comparable or better than previous fine-tuned models on out-of-domain data as reported in Appendix E.
+To assess variability, each test set is segmented into 10 chunks during evaluation resulting in a box-plot.
+Larger models perform better than previous fine-tuned models on out-of-domain data as reported in Appendix E.
 Notably, when answer-bearing contexts are provided (Fig.
-1d) versus when they are absent (Fig.
-1e), the models show significant performance gains, suggesting comprehension abilities in Sanskrit, a language characterized by high inflection.
+3d) versus when they are absent (Fig.
+3e), the models show significant performance gains, suggesting comprehension abilities in Sanskrit, a language characterized by high inflection.
 This behavior is however, not exhibited by smaller models when prompted in Sanskrit.
-3.1 Sanskrit vs English Prompts
-In our evaluations for san, we prompt in both English and Sanskrit, as mentioned earlier.
-Figure 1 shows that, for most models, prompting in English significantly outperforms prompting in Sanskrit—except for gpt-4o, where performance is largely similar.
-This provides partial evidence that these models are not instruction tuned on Sanskrit, aligning with trends observed by Muennighoff et al.
-(2023).
-In the case of gpt-4o, the comparable performance may indicate that the model has reached a saturation point to its best possible zero-shot performance as it is unlikely to have been trained on substantially more instruction tuning data than gpt-4o-mini.
-Figure 2: Confusion matrices from the NER task in san (a-d), lat (e-h) and grc (i-l), normalized across rows.
+3.1 Prompt Language: English versus Native
+During evaluation, we prompted models both in English and in each target language.
+As shown in Figure 3, English prompts generally outperform Sanskrit prompts, particularly with smaller models, providing partial evidence that these models have not been instruction-tuned on Sanskrit (Muennighoff et al., 2023).
+For Latin and Ancient Greek, this English-prompt advantage holds mainly for smaller models; larger models perform equally well, or even better, with native-language prompts (e.g., in Latin NER).
+This does not imply instruction tuning in these languages, since larger and smaller models likely saw comparable amounts of tuning data.
+Rather, it suggests that cross-lingual transfer is especially strong for Latin and Ancient Greek in larger models, reflecting their substantial influence on high-resource languages such as English.
+Figure 4: Confusion matrices from the NER task in san (a-d), lat (e-h) and grc (i-l), all with <en> prompts, normalized across rows.
 3.2 Misclassified Entities in NER
 Figure 4 displays confusion matrices for the NER task.
-Across all three languages, the smaller models exhibit more confusion among semantically related classes, while the larger models show fewer off-diagonal errors.
-In san, mythological entities like DEVA, ASURA, and RAKSHASA or semantically closed entities like KINGDOM vs CITY (eg., kośala vs ayodhyā) or FOREST (eg., daṇḍaka) vs GARDEN (eg., nandana, a heavenly garden)often get misclassified with each other in the smaller models.
+Across all three languages, the smaller models exhibit more confusion among semantically related classes (see G for descriptions of entity types), while the larger models show fewer off-diagonal errors.
+In san, mythological entities like Deva, Asura, and Rakshasa or semantically closed entities like Kingdom versus City (e.g., Kośala vs Ayodhyā) or Forest (e.g., Daṇḍaka) versus Garden (e.g., Nandana) often get misclassified with each other in the smaller models.
 For lat, entity type GRP proves troublesome for the smaller models, suggesting struggles in separating individual entities from collective ones.
-In grc, categories such as LOC and ORG show higher confusion counts akin to GRP in lat while confusion between GOD and PERSON seems similar to what we have just discussed in Sanskrit.
+In grc, categories such as LOC and ORG show higher confusion counts akin to GRP in lat while confusion between God and Persons seems similar to what was discussed for Sanskrit.
 In contrast, much clearer boundaries emerge in the larger models’ confusion matrices.
-In many of these cases, the domain or style of the texts—especially if they involve mythological or archaic terms typical of classical texts— can be understood to influence performance, as models with less exposure to specialized terminology may conflate related entity types.
-3.3 Inflection in Answers of san QA
+In many of these cases, the domain or style of the texts, especially if they involve mythological or archaic terms typical of classical texts, can be understood to influence performance, as models with less exposure to specialized terminology may conflate related entity types.
+3.3 Inflection in Answers in Sanskrit QA
 LLM Closed-book + RAG-BM25
@@
 Table 2: Comparison of EM scores in san QA (English prompts) when predicted and gold answers are considered with inflection or lemmatized.
-In the Sanskrit QA task, models are expected to generate single-word answers with correct inflection.
+In the Sanskrit question-answering task, models are expected to generate single-word answers with the correct inflection.
 For computing exact match (EM) scores, we manually identified all acceptable answers, excluding those with incorrect inflection (e.g., wrong case or gender endings).
-To quantify inflection errors, we also computed EM scores on lemmatized versions of both the gold-standard and predicted answers, as shown in Table 2.
-Most models exhibit only a slight increase in EM scores on lemmatized answers, suggesting that inflection errors are relatively minor—a finding corroborated by manual inspection.
+To quantify inflection errors, we also calculated EM scores on lemmatized versions of the gold standard and predicted answers, as shown in Table 2.
+Most models show only a slight increase in EM scores on lemmatized answers, suggesting that inflection errors are relatively minor, a finding corroborated by manual inspection.
 Future work could extend this analysis to investigate inflection accuracy in full sentence generation within broader natural language generation scenarios.
-3.4 Knowledge-Graph QA
-Additionally, we explored the use of incomplete knowledge graphs (KGs) for Sanskrit QA.
+3.4 Sanskrit Orthography: Devanagari versus IAST
+LLM MT (BLEU) NER (Macro F1-BI)
+Devanagari IAST Devanagari IAST
+gpt-4o 0.179 0.165 0.637 0.599
+llama-v3p1-405b-instruct 0.193 0.148 0.561 0.556
+gpt-4o-mini 0.135 0.099 0.359 0.318
+llama-v3p1-8b-instruct 0.120 0.063 0.164 0.149
+Table 3: Comparison of performances in san MT and NER (English prompts) when the input sentences are given Devanagari script or in IAST script.
+So far, we have shown robust cross-lingual generalization in the models.
+We now turn to one possible underlying mechanism—orthographic transfer—where models benefit from shared scripts across languages.
+Prior work has identified orthography as a key factor in cross-lingual transfer for LLMs (Muller et al., 2021; Fujinuma et al., 2022).
+To isolate this effect, we re-ran our Sanskrit NER and MT experiments (using English prompts) in Roman-based IAST transliteration instead of Devanagari.
+Table 3 compares performance in both scripts.
+Models perform better with the Devanagari script, which is shared by higher-resource relatives like Hindi and Marathi, reinforcing the importance of script sharing.
+However, results in IAST are only slightly lower, suggesting that Roman-based transliterations also feature prominently in the pre-training data.
+In future, we will investigate whether model outputs are consistent across both scripts, that is, whether these LLMs are effectively digraphic in Sanskrit.
+3.5 Knowledge-Graph Question-Answering
+Additionally, we explore the use of knowledge graphs (KGs) for Sanskrit QA.
 We evaluated a KG derived from the Bhāvaprakāśanighaṇṭu text (Terdalkar et al., 2023) and constructed a small KG for Rāmāyaṇa (details in Appendix F).
-Using the Think-On-Graph (ToG) paradigm (Sun et al., 2024), which iteratively explores KG paths for answer retrieval in a zero-shot, training-free manner (Xu et al., 2024), we observed that gpt-4o could effectively execute this method.
-Although it occasionally extracted correct answers, its performance did not significantly exceed that of the closed-book setting, possibly due to the incompleteness of the KGs.
+Using the Think-On-Graph (ToG) paradigm (Sun et al., 2024), which iteratively explores the KG paths for answer retrieval in a training-free zero-shot manner (Xu et al., 2024), we observed that gpt-4o could effectively execute this method.
+Although it occasionally extracted correct answers, its performance did not significantly exceed that of the closed-book setting, most likely due to the incompleteness of the KGs (see §F.3).
 Future work may focus on developing more comprehensive KGs to enhance this retrieval method.
 4 Conclusions
-In summary, our zero-shot evaluations demonstrate that larger language models exhibit robust cross-lingual generalization across diverse natural language understanding tasks in classical languages—including NER, machine translation, and QA.
-Notably, the significant performance gains achieved when answer-bearing contexts are provided—particularly in Sanskrit QA—suggest comprehension abilities in highly inflected languages.
+In summary, our zero-shot evaluations demonstrate that larger language models exhibit robust cross-lingual generalization across diverse natural language understanding tasks in classical languages, including NER, machine translation, and QA.
+Notably, the significant performance gains achieved when answer-bearing contexts are provided, particularly in Sanskrit QA, suggest comprehension abilities in highly inflected languages.
 Moreover, our contribution of a novel Sanskrit QA dataset provides a valuable resource for evaluating and benchmarking LLM performance on classical language tasks.
 Importantly, these models have not been explicitly instruction tuned on Sanskrit, Latin, or Ancient Greek—evidenced by the superior performance achieved when using English prompts for Sanskrit—which indicates that their zero-shot performance is attributable solely to cross-lingual generalization.
-Future work should focus on expanding dataset coverage and exploring additional classical languages and tasks, further advancing our understanding of cross-lingual generalization in LLMs and its applications in digital humanities and multilingual NLP research.
+Future work will focus on expanding dataset coverage, knowledge graphs and exploring additional classical languages and tasks, further advancing our understanding of cross-lingual generalization in LLMs and its applications in digital humanities and multilingual NLP research.
+Acknowledgements
+This research is financially supported by the Indian Knowledge Systems (IKS) Division of Ministry of Education, Govt.
+of India (project number AICTE/IKS/RFP1/2021-22/12).
+Mahesh Akavarapu received funding from Volkswagen Foundation under the Phylomilia project within the Pioneering Projects funding line.
+We also thank anonymous reviewers and the Area Chairs for their comments that have helped improve the paper.
 Limitations
@@
 Our evaluation relies exclusively on zero-shot performance, as the models have not been explicitly instruction tuned on these languages; this design choice may obscure potential benefits achievable through targeted fine-tuning.
+Further, a few datasets we experimented were released within the models’ knowledge cut-off dates raising the issue of data contamination.
+Among these, only Ancient Greek MT exhibits anomalously high performance, suggesting possible exposure.
+In general, NER, owing to its structural data should be less susceptible to contamination than MT.
 Furthermore, the effectiveness of our BM25-based retrieval approach depends heavily on preprocessing steps such as lemmatization, which might not optimally address all linguistic variations in highly inflected languages.
@@
 Ethics Statement
-Classical Sanskrit epics hold deep cultural and religious significance in Indian traditions, and similarly, Āyurveda represents a sensitive area within healthcare.
+Classical Sanskrit epics hold deep cultural and religious significance in Indian traditions, and similarly, Āyurveda represents a revered tradition-bound area within healthcare.
 We acknowledge that any research involving these subjects must be conducted with particular care.
@@
 Although our experiments indicate that paradigms like RAG produce more grounded and, hence, potentially safer outputs, there is no assurance that the responses from current LLMs in these domains meet clinical or religious safety standards.
-Accordingly, the authors do not endorse using the datasets beyond the scope of linguistic research.
-These datasets will be released under an open-source, non-commercial license, and all annotators have been compensated at fair, standard rates.
-Appendix A Question Answering Dataset
+Consequently, the authors do not endorse using the datasets beyond the scope of linguistic research.
+These datasets are released for open-source, non-commercial use, and all annotators have been compensated at fair, standard rates.
+Appendix
+Appendix A Prompts
+The Sanskrit prompts are in Devanagari script. In this appendix, we provide these prompts transliterated in IAST scheme.
+A.1 Prompts for Named Entity Recognition
+Prompt in English
+Recognize the named entities from the following sentence in {LANGUAGE}. The valid tags are {ENTITY TYPES}. Do not provide explanation and do not list out entries of ‘O’. Example:
+Sentence: <word_1> <word_2> <word_3> <word_4> <word_5>
+Output: {{‘B-<entity1>’: [‘<word_1>’, ‘<word_4>’], ‘B-<entity2>’:[‘<word_5>’]}}
+Sentence: {INPUT}
+Output:
+(The example is never a real sentence and is only provided to specify the output structure. Hence, the evaluations are strictly zero-shot.)
+Prompt in Sanskrit
+adho datta vākye nāmakṛtāḥ sattvāḥ (named entities) abhijānīhi. tadapi vivṛtam mā kuru, kevalam pṛṣṭa viṣayasya uttaram dehi. api ca ‘O’-sambandhitāni na deyāni.
+sattvāḥ eteṣu vargeṣu vartante - {ENTITY TYPES}. udāharaṇāya -
+vākyam: <padam_1> <padam_2> <padam_3> <padam_4> <padam_5>
+phalitam: {{ ‘B-<sattvaḥ1>’: [‘<padam_1>’, ‘<padam_4>’], ‘B-<sattvaḥ2>’: [‘<padam_5>’]}}
+vākyam: {INPUT}
+phalitam:
+Prompt in Latin
+Agnosce nomina propria (named entities) ex hac sententia Latina. Notae validae sunt {ENTITY TYPES}. Explanationem ne praebeas nec elementa ‘O’ elenca. Exemplar:
+Sententia: <verbum_1> <verbum_2> <verbum_3> <verbum_4> <verbum_5>
+Productus: {{‘B-<entitatem1>’: [‘<verbum_1>’, ‘<verbum_4>’], ‘I-<entitatem1>’: [‘<verbum_2>’], ‘B-<entitatem3>’:[‘<verbum_5>’]}}
+Sententia: {INPUT}
+Productus:
+Prompt in Ancient Greek
+[…]
+[…]
+[…]
+$\pi\rho\acute{o}\tau\alpha\sigma\iota\varsigma$: <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_1> <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_2> <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_3> <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_4> <$\lambda\acute{\varepsilon}\xi\iota\varsigma$_5>
+[…]
+$\pi\rho\acute{o}\tau\alpha\sigma\iota\varsigma$: {INPUT}
+$\pi\alpha\rho\alpha\gamma\omega\gamma\acute{\eta}$:
+A.2 Prompts for Machine Translation
+Prompt in English
+Translate the following sentence in {LANGUAGE} into English. Do not give any explanations.
+Prompt in Sanskrit
+adho datta-saṃskṛta-vākyam āṅgle anuvādaya, tad api vivṛtam mā kuru -
+Prompt in Latin
+Verte hanc sententiam Latinam in Anglicam. Nullam explicationem praebe.
+Prompt in Ancient Greek
+[…]
+(Sanskrit QA Prompts)
+In the following prompts, TOPIC is either ‘Rāmāyaṇa’ or ‘Āyurveda’.
+A.3 Prompts for Closed-book QA
+Prompt in English
+Answer the question related to {TOPIC} in the Sanskrit only.
+Give a single word answer if reasoning is not demanded in the answer.
+With regards to how-questions, answer in a short phrase, there is no single word restriction.
+{QUESTION} {CHOICES}
+Prompt in Sanskrit
+tvayā saṃskṛta-bhāṣāyām eva vaktavyam.
+na tu anyāsu bhāṣāsu.
+adhaḥ {TOPIC}-sambandhe pṛṣṭa-praśnasya pratyuttaraṃ dehi.
+tadapi ekenaiva padena yadi uttare kāraṇam nāpekṣitam.
+katham kimartham ityādiṣu ekena laghu vākyena uttaraṃ dehi atra tu eka-pada-niyamaḥ nāsti.
+{QUESTION} {CHOICES}
+A.4 Prompts for RAG-QA
+Prompt in English
+Answer the following question related to {TOPIC} in Sanskrit only.
+Give a single word answer if reasoning is not demanded in the answer.
+With regards to how-questions, answer in a short phrase.
+Also take help from the contexts provided.
+The contexts may not always be relevant."
+contexts: {CONTEXTS}
+question:{QUESTION} {CHOICES}
+Prompt in Sanskrit
+tvayā saṃskṛta-bhāṣāyām eva vaktavyam.
+na tu anyāsu bhāṣāsu.
+adhaḥ {TOPIC}-sambandhe pṛṣṭa-praśnasya pratyuttaraṃ dehi.
+tadapi ekenaiva padena, yāvad laghu śakyaṃ tāvad, taṃ punaḥ vivṛtam mā kuru.
+api ca yathā’vaśyam adhaḥ datta-sandarbhebhyaḥ ekatamāt sahāyyaṃ gṛhāṇa.
+tattu sarvadā sādhu iti nā’sti pratītiḥ.
+sandarbhāḥ: {CONTEXTS}
+praśnaḥ: {QUESTION} {CHOICES}
+Appendix B Question Answering Dataset
 In this appendix, we describe the creation of Sanskrit QA dataset.
@@
 For such questions (only about 50), the answers can be paraphrased variously, thereby requiring manual evaluation.
-The broad semantic and domain-specific categories of the questions are detailed in Tables 4 and 4.
+The broad semantic and domain-specific categories of the questions are detailed in Tables 5 and 5.
 Category Description #Q
@@
 Table 5: Question Categories for Āyurveda QA Dataset
-Appendix B Retrieval Augmented Generation
+Appendix C Retrieval Augmented Generation
 In the RAG paradigm, the LLM is provided with additional context that consists of top-$k$ passages retrieved from the original texts.
@@
 As a result, the lemmatizer trained on Classical Sanskrit and thereby, our entire retrieval pipeline may not be appropriate on such corpora and hence are not comparable.
-We evaluated RAG with different $k$ using GPT-4o with prompts in Sanskrit.
-The EM scores are plotted in Fig.
-3.
-It can be seen that BM25 outperforms AvgFT and AvgGV, it may also be inferred from both plots that $k=4$ would be optimal.
-At this value of $k$, the performance of RAG in the case of Bhāvaprakāśanighaṇṭu already starts to deteriorate, although not significantly.
-Keeping both plots in mind, we have fixed the value of $k$ at 4.
-Figure 3: Effect of $k$ on RAG, denoting the number of top best matching text chunks retrieved, on the performances of GPT-4o with retrievers based on BM25, averaged FastText (AvgFT) and GloVe (AvgGV) embeddings respectively of datasets Rāmāyaṇa (left) and Bhāvaprakāśanighaṇṭu (right).
-The prompts for RAG are detailed in Appx. C.4.
-Appendix C Prompts
-The Sanskrit prompts are in Devanāgari script.
-In this appendix, we provide the prompts (transliterated in IAST scheme [^5] https://en.wikipedia.org/wiki/International_Alphabet_of_Sanskrit_Transliteration).
-C.1 Prompts for Named Entity Recognition
-Prompt in English
-Recognize the named entities from the following sentence in {LANGUAGE}. The valid tags are {ENTITY TYPES}. Do not provide explanation and do not list out entries of ‘O’.
-Example:{EXAMPLE}
-(Note: EXAMPLE in Sanskrit is:
-daśarathaḥ ayodhyāyāḥ rājā āsīt
-{‘B-HUMAN’: [‘daśarathaḥ’], ‘B-CITY’: [‘ayodhyāyāḥ’]}
-An example is provided to make the output structure clear.)
-Prompt in Sanskrit
-adho datta vākye named entities abhijānīhi (NER kuru). tadapi vivṛtam mā kuru, kevalam pṛṣṭa viṣayasya uttaram dehi.
-entities eteṣu vargeṣu vartante - {ENTITY TYPES}
-udāharaṇāya - {EXAMPLE}
-C.2 Prompts for Machine Translation
-Prompt in English
-Translate the following sentence in {LANGUAGE} into English. Do not give any explanations.
-Prompt in Sanskrit
-adho datta-saṃskṛta-vākyam āṅgle anuvādaya, tad api vivṛtam mā kuru -
-(Sanskrit QA Prompts)
-In the following prompts, TOPIC is either ‘Rāmāyaṇa’ or ‘Āyurveda’.
-C.3 Prompts for Closed-book QA
-Prompt in English
-Answer the question related to {TOPIC} in the Sanskrit only.
-Give a single word answer if reasoning is not demanded in the answer.
-With regards to how-questions, answer in a short phrase, there is no single word restriction.
-{QUESTION} {CHOICES}
-Prompt in Sanskrit
-tvayā saṃskṛta-bhāṣāyām eva vaktavyam.
-na tu anyāsu bhāṣāsu.
-adhaḥ {TOPIC}-sambandhe pṛṣṭa-praśnasya pratyuttaraṃ dehi.
-tadapi ekenaiva padena yadi uttare kāraṇam nāpekṣitam.
-katham kimartham ityādiṣu ekena laghu vākyena uttaraṃ dehi atra tu eka-pada-niyamaḥ nāsti.
-{QUESTION} {CHOICES}
-C.4 Prompts for RAG-QA
-Prompt in English
-Answer the following question related to {TOPIC} in Sanskrit only.
-Give a single word answer if reasoning is not demanded in the answer.
-With regards to how-questions, answer in a short phrase.
-Also take help from the contexts provided.
-The contexts may not always be relevant."
-contexts: {CONTEXTS}
-question:{QUESTION} {CHOICES}
-Prompt in Sanskrit
-tvayā saṃskṛta-bhāṣāyām eva vaktavyam.
-na tu anyāsu bhāṣāsu.
-adhaḥ {TOPIC}-sambandhe pṛṣṭa-praśnasya pratyuttaraṃ dehi.
-tadapi ekenaiva padena, yāvad laghu śakyaṃ tāvad, taṃ punaḥ vivṛtam mā kuru.
-api ca yathā’vaśyam adhaḥ datta-sandarbhebhyaḥ ekatamāt sahāyyaṃ gṛhāṇa.
-tattu sarvadā sādhu iti nā’sti pratītiḥ.
-sandarbhāḥ: {CONTEXTS}
-praśnaḥ: {QUESTION} {CHOICES}
+The prompts for RAG are detailed in Appx. A.4.
 Appendix D Implementation
@@
 llama-3.1-405b-instruct 17.1
-MT (san-eng)
+MT (san-eng) on Mann ki Baat dataset
 Model Macro F1 (BI)
@@
 LatinBERT2 (Beersmans et al., 2023) 0.50
-gpt-4o 0.51
-llama-3.1-405b-instruct 0.49
-NER (lat)
+gpt-4o 0.55
+llama-3.1-405b-instruct 0.36
+NER (lat) on Ars Amatoria dataset
 Table 6: Comparison of out of domain performances of LLMs against previously reported fine-tuned models.
 In Table 6, we compare the out-of-domain performance of our evaluated models against previously reported fine-tuned models.
-For MT (san-eng), open-source model llama-3.1-405b-instruct achieves a BLEU score of 17.1, outperforming both Google Trans and IndicTrans, while for NER (lat), the performance of LLMs is roughly comparable to that of fine-tuned LatinBERT variants.
-Although fine-tuned models yield superior results on in-domain data, our findings indicate that multilingual LLMs are competitive—or even superior—in their zero-shot generalization.
+For MT (san-eng) on Mann ki Baat dataset (Maheshwari et al., 2024), open-source model llama-3.1-405b-instruct outperforms both Google Trans and IndicTrans, while for NER (lat) on Ovid’s Ars Amatoria dataset (Beersmans et al., 2023), the performance of gpt-4o is better than that of fine-tuned LatinBERT variants.
+Although fine-tuned models yield superior results on in-domain data, our findings indicate that multilingual LLMs are superior in their zero-shot generalization.
 Appendix F LLMs with Knowledge Graphs
-Arriving at an answer by an LLM integrated with a knowledge graph (KG) through Think-on-Graph (ToG) (Sun et al., 2024) paradigm involves several prompting steps for each hop from starting entity nodes.
+Figure 5: Overview of augmenting a LLM with a knowledge graph (KG) through Think-on-Graph (ToG) paradigm.
+Arriving at an answer by an LLM integrated with a knowledge graph (KG) through Think-on-Graph (ToG) (Sun et al., 2024) paradigm involves several prompting steps for each hop from starting entity nodes as illustrated in Fig.
+5.
 Firstly, the LLM lists entities from the input questions further lemmatized by our lemmatizer previously described.
@@
 katham kimartham ityādiṣu ekena laghu vākyena uttaraṃ dehi atra tu eka-pada-niyamaḥ nāsti.
-api ca yathā’vaśyam adhaḥ dattaiḥ knowledge-graph-taḥ niṣkarṣita-viṣayaiḥ sahạyyaṃ gṛhāṇa. tattu sarvadā sādhu iti nā’sti pratītiḥ. uttaram yāvad laghu śakyam tāvat laghu bhavet.
+api ca yathā’vaśyam adhaḥ dattaiḥ knowledge-graph-taḥ niṣkarṣita-viṣayaiḥ sahāyyaṃ gṛhāṇa. tattu sarvadā sādhu iti nā’sti pratītiḥ. uttaram yāvad laghu śakyam tāvat laghu bhavet.
 human praśnaḥ: {QUESTION} {CHOICES}
@@
 uttaram:
+F.3 LLM-KG Results
+Method gpt-4o claude-3.5-sonnet gemini-1.5-pro mistral-large-2 llama-3.1-405b-instruct
+Closed-book 0.381 0.242 0.148 0.333 0.346
+RAG-BM25 0.478 0.521 0.459 0.434 0.323
+LLM-KG 0.381 0.254 - 0.341 -
+Table 7: Exact Match (Scores) of various models (including those not part of main experiments) in Sanskrit Question-Answering task (Sanskrit Prompts) with LLM-KG paradigm compared against zero-shot and RAG-BM25 paradigms.
+The LLM-KG paradigm was evaluated exclusively using Sanskrit prompts on the two QA datasets and included additional models not part of the main experiments—namely, claude-3.5-sonnet (AnthropicAI, 2024), gemini-1.5-pro (Google, 2024), and mistral-large-2 (MistralAI, 2024).
+Table 7 presents the results in comparison with the closed-book and RAG-BM25 paradigms.
+Overall, performance gains from closed-book to LLM-KG are modest and fall short of the improvements observed with RAG.
+This may be partly attributed to the complexity of the LLM-KG setup, which requires multi-step prompting and adherence to a structured output format.
+Notably, models like gemini-1.5-pro and llama-3.1 frequently fail to follow this structured format, rendering them ineffective for running ToG.
+The strict formatting requirements may also pose challenges for other models, particularly those less adapted to Sanskrit.
+Interestingly, while claude-3.5-sonnet achieves the best results with RAG-BM25, it lags behind gpt-4o and mistral-large-2 in both the closed-book and LLM-KG paradigms.
+Method gpt-4o claude-3.5-sonnet mistral-large-2
+closed-book 0.32 0.21 0.25
+LLM-KG 0.34 0.34 0.35
+Method gpt-4o claude-3.5-sonnet mistral-large-2
+closed-book 0.40 0.25 0.36
+LLM-KG 0.39 0.23 0.34
+Table 8: Comparison of Exact Match (EM) scores between closed-book and LLM-KG paradigms for selected questions when the answer (a) can likely be inferred from KG and (b) cannot be inferred from KG.
+Table 8 presents a breakdown of performance based on whether the question topics are covered in the current KG—specifically, the kingdoms category (27 questions) in the Rāmāyaṇa dataset and the annotated chapters (299 questions) in Bhāvaprakāśanighaṇṭu.
+For these subsets, which are likely answerable from the KG, LLM-KG shows clear improvements over the closed-book setting, indicating that access to a near-complete KG can significantly enhance performance.
+In contrast, for questions outside these categories or chapters, no such improvement is observed, reinforcing the hypothesis that KG completeness is crucial for the effectiveness of LLM-KG.
+Determining domains where knowledge graphs may outperform or be more appropriate than RAG remains an open question for future research.
+Appendix G Categories for Named Entity Recognition
+The categories for NER in Sanskrit, Ancient Greek, and Latin, along with their rough translation and brief explanations, wherever applicable, are provided here.
+Entity Type Translation Description
+Manuṣya Human A mortal human being
+Deva Deity Divine celestial being; god or goddess
+Gandharva $\sim$ Heavenly musician in the service of the gods
+Apsaras $\sim$ Beautiful female spirits known for dance and charm
+Yakṣa $\sim$ Guardian spirit of natural treasures.
+Kinnara $\sim$ Certain Semi-divine beings
+Rākṣasa $\sim$ Malevolent being
+Asura Anti-god Powerful beings opposed to the gods
+Vānara Monkey-being Monkey-like humanoid
+Bhallūka Bear-being Bear or Bear-like humanoid
+Gṛdhra Vulture-being Vulture-like being
+Ṛkṣa Bear-being Bear-like humanoid
+Garuḍa Eagle-being Eagle-like being
+Nāga Serpent-being Semi-divine serpent race
+Svarga Heaven Abode of the gods
+Naraka Hell Realm of punishment after death
+Nadī River Flowing body of freshwater
+Sāgara Sea Vast saltwater body
+Sarovara Lake Large inland water body
+Kūpa Well Man-made water source
+Tīra Riverbank Edge or shore of a river
+Dvīpa Island Land surrounded by water
+Parvata Mountain Large natural elevation of earth
+Nagara City Urban settlement or metropolis
+Tīrtha Sacred Place Holy pilgrimage spot, often near water
+Grāma Village Small rural settlement
+Rājya Kingdom Territory ruled by a king
+Vana Forest Dense growth of trees; wilderness
+Udyāna Garden Cultivated green space
+Marubhūmi Desert Dry, arid region
+Prāsāda Palace Royal residence
+Mandira Temple Sacred structure for worship
+Āśrama Hermitage Secluded place for spiritual practice
+Gṛha House Dwelling or home
+Kuṭīra Hut Small and simple shelter
+Guhā Cave Natural underground chamber
+Mārga Road Pathway or route
+Ratha Chariot Two- or four-wheeled ancient vehicle
+Vimāna Airborne Vehicle Flying chariot or aircraft
+Khadga Sword Bladed weapon
+Dhanus Bow Weapon for shooting arrows
+Bāṇa Arrow Projectile shot from a bow
+Cakra Discus Spinning circular weapon
+Gadā Mace Blunt weapon, often spiked
+Tomara Javelin Thrown spear or missile
+Śūla Spear Long-shafted piercing weapon
+Kavaca Shield Defensive armor piece
+Kañcuka Armor Protective body gear
+Paraśu Axe Bladed tool/weapon
+Astra Divine Weapon Supernatural weapon, often invoked
+Ābharaṇa Ornament Decorative jewelry
+Śaṅkha Conch Sacred spiral shell
+Vādya Musical Instrument Instrument used in music
+Nāṇa Currency Form of money or coin
+Kula Clan Extended family or lineage
+Jāti Species Species/Socio-economical Group
+Gaṇa Tribe / Group Assembly or community
+Ṛtu Season Climatic period of the year
+Saṃvatsara Year Vedic year cycle
+Māsa Month Lunar or solar month
+Tithi Lunar Day Phase in the moon’s waxing/waning
+Pakṣa Fortnight Half of a lunar month
+Ayana Solstice Cycle Six-month movement of the sun
+Yuga Epoch Cosmic age or era
+Yoga Astronomical Combination Planetary conjunction
+Karaṇa Half of Tithi Subdivision of a lunar day
+Muhūrta Moment / Auspicious Time Small unit of time (about 48 minutes)
+Lagna Ascendant Zodiac rising at time of birth
+Graha Planet Celestial influencer
+Nakṣatra Lunar Mansion One of 27 lunar constellations
+Rāśi Zodiac Sign Segment of the zodiac
+Dhuma-ketu Comet Celestial object with a tail
+Utsava Festival Celebratory event
+Pūjā Worship Ritual offering and prayer
+Yajña Vedic Sacrifice Sacred fire ritual
+Upacāra Ritual Offering Ceremonial gesture or item
+Saṃskāra Life-Cycle Rite Hindu ritual of life transition
+Aniścita Undecided Something that is not yet determined
+Vṛkṣa Tree Large woody plant
+Guccha Shrub Small bushy plant
+Lata Vine Climbing or trailing plant
+Puṣpa Flower Blossom of a plant
+Phala Fruit Edible plant product
+Patra Leaf Green foliage part
+Stambha Stem Main structural plant part
+Tvak Bark Outer layer of tree
+Mūla Root Underground part of plant
+Pakṣī Bird Feathered flying animal
+Sarpa Snake Legless reptile
+Table 9: Entity types occuring in Sanskrit NER
+Entity Type Description
+NORP Ethnic groups, demonyms, schools
+ORG Organizations
+GOD Supernatural beings
+LANGUAGE Languages and dialects
+LOC Cities, empires, rivers, mountains, and so forth.
+PERSON Individual persons
+Table 10: Entity types occuring in Ancient Greek NER (Myerston, 2025). The types without descriptions—EVENT and WORK—have very few occurances in the dataset.
+Entity Type Description
+PER Person
+LOC Locations, places
+GRP Other groups such as tribes
+Table 11: Entity types occuring in Latin NER are quite standard types.